OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)? y - 2 = 0 x - 2 = 0 y - x - 2 = 0

OpenStudy (anonymous):

that's y=2.let me explain it

OpenStudy (anonymous):

.points (-1, 2) and (5, 2) slope m=(y2-y1)/(x2-x1) m=(2-2)/(-1-5)=0

OpenStudy (anonymous):

y-y1=m(x-x1) y-2=0(x-5) y=2

OpenStudy (anonymous):

you have these point : x1= -1 , y1=2 and x2=5 and y2=2.... and we know that a line equation in general is : y-y1 = m (x-x1) which m is slope of the line , and x1,y1 is one of the points (e.g first point) and m = (y2-y1)/(x2-x1). done! :)

OpenStudy (anonymous):

exactly as mark_o mentioned.